### v1

# HF radar estimation of ocean wave parameters: second-order Doppler spectrum versus Bragg wave modulation approach

Abstract
We propose an original technique for the HF radar estimation of the main sea state parameters by exploiting the amplitude modulation of the radar signal time series. While the classical method for ocean wave measurement is based on the second-order ocean Doppler spectrum, this alternative approach makes use of the dominant first-order Bragg echo in the radar signal, which has a larger signal-to-noise ratio and allows for a farther spatial coverage. We apply this method to a yearly set of HF radar data in the vicinity of Toulon (Mediterranean coast of France) and compare it with the classical second-order method, where we use a nearby buoy record as ground truth. The Doppler spectrum-based method is slightly more accurate in estimating the significant wave height and the mean period at a short distance but the temporal approach allows for a significant increase in range and better captures the variability of wave parameters over time.
journal: Remote Sensing of Environment
[…]
[1]organization=MIO (Univ Toulon, Aix-Marseille Univ, CNRS, IRD),city=Toulon, country=France
1 Introduction
High-frequency radars (HFR) have been used for about half a century for the monitoring of coastal surface currents (Barrick et al., (1977)). As a secondary geophysical product, HFR also provide an indirect measurement of sea state parameters and wave spectra (e.g. Wyatt, (2021) for a recent review). However, this type of estimation is less accurate and less straightforward than the surface current estimation, for both theoretical and practical reasons. On the theoretical level, while the radial surface current only requires the measurement of the dominant first-order Bragg peaks, the wave spectrum is related to the side-band continuous components of the ocean Doppler spectrum surrounding the Bragg peaks. This second-order contribution has a complex analytical expression in the form of a nonlinear integral equation involving a coupling coefficient and quadratic combinations of the wave spectra at pairs of frequencies satisfying a resonance condition Hasselmann, (1971); Barrick, (1972); Lipa and Barrick, (1986). Under some restrictive assumptions, Barrick, 1977b ; Barrick, 1977a derived in $1977$ a method to extract the main sea state parameters from this second-order integral and accordingly, from the measured ocean Doppler spectrum. He showed that the significant wave height ($H_{s}$) and the peak wave period ($T_{p}$) can be expressed as simple ratios of weighted integrals involving the first- and second-order Doppler spectra. Soon after, linear and nonlinear inversion methods for retrieving the ocean wave spectrum from one single radar site were developed in a series of papers (e.g. Lipa, (1977); Wyatt, (1990); Hisaki, (1996); Wyatt, (2021)). However, the success of these methods for inverting the sea state parameters or the wave spectrum is bound to the ability to measure the second-order Doppler spectrum with a sufficient signal-to-noise ratio. Except for the high sea state and the highest radar frequencies, the second-order Doppler spectrum is usually several tens of dB below the level of the first-order peaks and reaches sooner the noise threshold when the range is increased. Hence, wave estimation methods based on the second-order Doppler analysis are confined to a restricted radar coverage in the near to middle range, depending on the radar frequency. Another issue is the clear identification of the second-order contribution and its separation from the first-order Bragg components, which can be ambiguous especially in the case of strong currents or in the absence of azimuthal resolution. For these reasons, empirical methods using the second-order Doppler spectrum (e.g. Essen et al., (1999); Gurgel et al., (2006); Lopez et al., (2016)) or empirical adjustments of Barrick’s method Heron and Heron, (1998); Maresca Jr and Georges, (1980); Ramos et al., (2009); Alattabi et al., (2019); Al-Attabi et al., (2021); Basañez et al., (2020) are often preferred to the original approach of Barrick. Note that these methods necessitate the calculation of Doppler spectra that are resolved both in range and azimuth and can thus be associated with a reduced sea surface patch with constant sea state parameters. In the case of compact antenna arrays, the Doppler spectra are omnidirectional, that is, they cannot be resolved in azimuth. This does not allow for the measurement of the wave frequency spectrum from a single station. Nevertheless, the $H_{s}$ can be estimated on range arcs adapting Barrick’s original methods. The long-term performances of such measurements have been assessed by comparison with buoys and models and are quite variable depending on sea state, range, and direction of waves (see e.g. Roarty et al., (2015); Orasi et al., (2018); Saviano et al., (2020) or Lorente et al., (2022) for a recent review).
Figure 1: Location of HF radar coverage a) in the western Mediterranean Sea and b) zooming in on the area. The red dot marks the location of Fort Peyras, the blue dot that of the CANDHIS buoy, and the white lines represent the resolution in ranges (every $7.5km$) and azimuths (every $5\rm{{}^{o}}$) of the radar coverage. The orange line a40 marks the main azimuth and the orange arcs of circles mark the range numbers r8, r20, and r35 used for the HFR comparison. The background color is the bathymetry measured in $m$. Toponyms cited in the main text appear in the figure.
In this paper, we propose a novel approach to sea state estimation based on the analysis of the first-order Bragg component of the Doppler spectrum. We exploit the amplitude modulation of the received radar time series to infer the $H_{s}$ and the $T_{p}$. The idea of using the first-order backscattered signal to estimate the sea state parameters has been recently put forward in Shahidi and Gill, (2020); Shahidi et al., (2024) based on some mathematical approximations starting from the expression of the electric field backscattered from a patch of the sea surface. It has been found that the $H_{s}$ is simply proportional to the variance over time of the first-order received electric field and that the $T_{p}$ can be obtained from the peak frequency of the same times series after integration over range. The method has been assessed with limited but promising experimental tests. Our approach, which was developed independently, is based on simple physical heuristic arguments and reaches the same conclusion that the sea state parameters can be derived from the statistical properties of the radar signal time series (in which the first-order Bragg contribution is dominant) and does not require any Doppler processing nor the identification of the second-order contribution. Our formulation, however, is quite different in that it relies on the calculation of the analytical and anti-analytical radar signal, a technique that allows us to isolate the contribution of long waves in the backscattered field.
The paper serves a double objective. First, we test the performance of Barrick’s method based on a yearly data set of HF measurements with a radar located near Toulon. Its accuracy is very sensitive to some tuning parameters in its numerical implementation, which is part of the tacit know-how of the users and is rarely mentioned in the literature. A key step is the identification of the separation between the first- and second-order Doppler spectrum, for which we propose an empirical but explicit criterion. The second objective is to test the aforementioned new method based on the temporal properties of the backscattered signal. The HFR data set is described in section 2, together with in-situ measurements and numerical models which are used for the assessment. The radar processing methods are presented in Section 3 and a variety of statistical comparisons over one year of data is given in Section 4 while discussing the results obtained.
Figure 2: a) Mean value (in meter) of $H_{s}$ over the year 2020 in the radar coverage according to the WWIII model; b) Root-mean-square difference (in meter) between the $H_{s}$ at any location and the $H_{s}$ at the CANDHIS buoy location, according to the WWIII model. The black dashed line marks the azimuth 40 and the black circles mark the ranges r8, r12, and r35, respectively, used for the HFR comparison (see Figure 1).
2 Data
2.1 High Frequency Radar data
The Toulon HFR network, developed by WERA Helzel Messtechnik, consists of a combination of two transmitters and two receivers located at three different sites. On the island of Porquerolles, $27$ km southeast of Toulon, there is a stand-alone transmitter with a single omnidirectional antenna covering a wide area to the south. Cap Bénat, located $35$ km east of Toulon, hosts the first set of receivers, consisting of a linear array of $12$ active antennas oriented $70$ degrees counterclockwise from the north, with an antenna separation of $0.45\lambda$. The second set of transmitters and receivers is located at Fort Peyras, about $8$ kilometers southwest of Toulon, with a linear array of $12$ passive antennas oriented north-south, also with a spacing of $0.45\lambda$. This configuration of transmitters and receivers produces three bistatic pairs and one monostatic pair Guérin et al., (2019); Dumas et al., (2020); Dumas and Guérin, (2023).
In this study, we use both omnidirectional and directional monostatic signals collected by the Fort Peyras site (Figure 1) during $2020$, a year for which data availability was over $91\%$. This radar operates in a $16.15$ MHz center frequency band with an assigned bandwidth of $100$ kHz. Depending on weather conditions, it can cover a range of $60$ km with a range resolution of $1.5$ km and an angular sector of about $110\rm{{}^{o}}$. The data consist of the complex I/Q voltage recorded on the set of $12$ receive antennas and sorted in range after a preliminary Doppler processing in the fast time. The sampling rate in the slow time is $0.26$ seconds and the time series are acquired in continuous blocks of $4096$ values (lasting about $18$ min) every $20$ min, which corresponds to the standard “SORT file” output format of WERA HF radar. The data are further processed in azimuth by applying the classical Beam Forming algorithm to the $12$ channels of the antenna array. This leads to a complex discrete time series $s(t)$ for each range and azimuth and each $20$ min sequence. Not applying the Beam Forming technique results in an omnidirectional signal for each antenna channel, with which it is only possible to know from what distance the signal is being received but not from which bearing.
2.2 Wave data
For the year $2020$ the time series of $H_{s}$ and $T_{p}$ are collected by a CANDHIS (Centre d’Archiveage National de Données de Houle In-Situ) buoy located in the vicinity of Porquerolles island ($42\rm{{}^{o}}58.00^{\prime}$, $6\rm{{}^{o}}12.29^{\prime}$, see Fig. 1). The data buoy are provided by the CEREMA (Climat et Territoires de Demain) and they are available in Cerema, (2024). The time resolution of the buoy data is $30$ minutes. During the same year, the results of the Wavewatch III model (WWIII) were available over the study area on an unstructured grid (with a spatial resolution between $400$m near the coast and $25$km offshore) forced by $10$-m wind fields obtained from the non-hydrostatic mesoscale Weather Research and Forecast model (for more details consult Lira-Loarca et al.,, 2022). The WWIII model has an hourly time resolution.
While the ground truth in the vicinity of the radar area is limited to the single location of the CANDHIS buoy, it is possible to use the WWIII model to complement the analysis. The spatial variability of $H_{s}$ under the radar coverage has been evaluated by calculating the root mean square difference over the whole year 2020 between the modeled $H_{s}$ at any location in the Toulon area and the modeled $H_{s}$ at the CANDHIS buoy location (Figure 2). As can be seen, this difference ranges from about $10$ cm at short distances from the radar to about 40 cm at long distances. Accepting this small to moderate difference, we can use the unique buoy measurement point as ground truth for validating the wave parameters estimated with the HFR dataset.
3 Material and Methods
Two different methods are used to estimate the ocean waves, $H_{s}$ and $T_{p}$, from HF radar signals. The first is the classical Second Order Doppler Spectrum (SODS) method, first introduced by Barrick, 1977a ; Barrick, 1977b , while the second is an original method that uses only the temporal variability of the first order peak and will be referred to as the “First Order Bragg Peak” (FOBP) method.
3.1 Second Order Doppler Spectra Method
The classical method for estimating $H_{s}$ and $T_{p}$ from the HF sea echo originates from the seminal work of Barrick, 1977a ; Barrick, 1977b and involves straightforward calculations based on the ocean Doppler spectrum. The hourly range-azimuth resolved Doppler spectra are classically obtained by forming the power spectrum of the radar signal:
$\sigma(\omega)=\langle\lvert\hat{s}(\omega)\rvert^{2}\rangle,$ (1)
where $\hat{s}(\omega)$ is computed in the standard way on a discrete set of frequencies using a Fast Fourier Transform of the time series with a Hamming attenuation window and an integration time of about $2$ min. To reduce the noise variance, a further incoherent summation is applied by taking a median average of the sample spectra $\lvert\hat{s}(\omega)\rvert^{2}$ over $15$ overlapping intervals within the $18$ min duration of each available SORT file (3 per hour). This will result in an average Doppler spectrum with a bandwidth of $4$ Hz (corresponding to a sampling rate of $0.26$ sec) and a frequency resolution of $\Delta f=1/T\simeq 0.008$ Hz. An example of such a Doppler spectrum is shown in Figure 3, a) for the $8$ range cell ($12$ km) and azimuth $40$ (marked with orange lines in Figure 1, b) for March $8$, $2020$. The corresponding range Doppler spectrum along the same direction is shown in Figure 4. In this example, the strongest first-order Bragg line extends to about $90$ km, while the strongest second-order sideband is visible up to $65$ km away from the radar. The dashed white circles highlight a power grid artifact ($50$ Hz harmonics) that contaminates the low Doppler frequency content of the radar signal at some ranges; this electrical artifact is particularly pronounced at the Fort Peyras site due to the imperfect power supply at this location. It does not affect the estimation of the radial current from the first-order Bragg lines but must be treated with caution when dealing with the full ocean Doppler spectrum.
To determine $H_{s}$, the ratio of the weighted integral of the second-order Doppler spectrum to the energy associated with the first-order peaks is analyzed (see figure 3,a).
[…]
where $\sigma_{1}$ and $\sigma_{2}$ are the first and second order Doppler spectra, respectively; ${\cal W}$ is the weighting function defined in Barrick, 1977b , $k_{0}$ is the scalar radio wavenumber, and
$\omega_{B}=2\pi\sqrt{\frac{g}{\pi\lambda}}$ (3)
is the Bragg frequency and $g=9.81\ m.s^{-2}$ is the gravitational constant.
The $T_{p}$ can also be determined from a different ratio of integrals based solely on the second-order Doppler spectrum:
[…]
where the bounds of the integrals can be either 00 and $\omega_{B}$ or $\omega_{B}$ and +$\infty$+ ∞, assuming that the positive frequency sidebands are the strongest (otherwise one should take the symmetric frequency intervals on the negative side). In this study, the maximum value of the integral of both intervals (from 00 to $\omega_{B}$ and from $\omega_{B}$ to +$\infty$+ ∞) is chosen for each time. The main challenge of this method is to effectively discriminate between first and second-order spectra and to accurately distinguish signal from noise. To overcome this, specific techniques have been developed that allow the precise setting of separation thresholds and the establishment of control criteria for integration regions. The effectiveness of this technique is limited by its dependence on the second-order Doppler spectrum, which restricts its application to shorter distances where this spectrum remains well above the noise level.
To determine the exact point of transition from first to second-order spectra, and thus avoid false boundaries, the following method is used: positive peaks (marked with a blue circle in Figure 3, b) and negative peaks (marked with red circles in Figure 3, b) are selected. The minimum negative peak (marked with a solid red circle in Figure 3, b) and the maximum positive peak at frequencies higher than the minimum negative peak (marked with a solid blue circle in Figure 3, b) are identified. We impose the condition that the intensity difference between the first-order peak and the minimum negative peak (d1) must be at least twice the difference between the minimum negative peak and the maximum positive peak (d2). If this condition is not met, the selected minimum negative peak is discarded and the process is repeated with another negative peak initially identified. At the same moment, the noise level is determined as the average intensity for frequencies higher than $1.75$ Hz (see Figure 3, a).
Figure 3: a) Doppler spectrum for March $8$, 2022022022020 at 00h at azimuth $40$ and range $8$ (see Figure 1, b). The dashed blue lines mark the first-order peaks. The second-order peaks are also marked in the figure, selecting the separation points between the first and second order with filled red dots. The noise threshold is marked with a dashed line and yellow shading. b) Zoom in on the right side of the positive first-order peak of the Doppler spectrum. The minimum/maximum peaks (red/blue points) when delimiting the first order from the second order of the spectrum are shown.
Figure 4: Range Doppler spectrum (in dB) for March $8$, 2022022022020 at 00h at azimuth $40$ (see Figure 1, b). The dashed red lines mark the Bragg frequency and the dashed white circles mark the few ranges where the zero-Doppler region is corrupted by power grid artifacts ($\sim 42$ and $82$ km).
3.2 First-order Bragg peak method
To achieve wave estimation over longer distances, a new method is proposed where working with the second-order spectrum becomes impossible. Instead, we exploit the modulation of the first-order Bragg signal by long waves, defined as waves significantly longer than the Bragg wave ($\omega_{p}<<\omega_{B}$). In a two-scale image, Bragg waves ride on long waves whose orbital motion modulates their phase velocity and amplitude. This process comes into play through a non-linear interaction between short and long waves, which heuristically involves the product of these two components. Assuming that the first-order Bragg scattering mechanism is dominant in the temporal radar signal $s(t)$, and considering both the approaching (positive frequencies) and receding (negative frequencies) components of the long waves and the Bragg waves, we can postulate the following simplified form for the received complex time series:
$s(t)\sim(R^{+}e^{i\omega_{p}t}+R^{-}e^{-i\omega_{p}t})(a^{+}e^{i\omega_{B}t}+a^{-}e^{-i\omega_{B}t})e^{i\omega_{u}t}$ (5)
where $t$ is time; $R^{\pm}$ and $a^{\pm}$ are the complex amplitudes with random phase of the long and Bragg waves respectively, $\omega_{p}$ is the peak frequency of the long wave and $\omega_{u}=4\pi u/\lambda$ is the additional frequency shift due to the radial surface current $u$. We expect $H_{s}$ to be proportional to the variance of the long wave modulation, i.e:
$H_{s}\sim\mathrm{Var}{(R^{+})}+\mathrm{Var}{(R^{-})}$ (6)
Note that the oscillation factor due to long wave modulation “survives” the radar cell averaging process because the wavelength of the longest waves is not negligible compared to the radar range resolution. For typical scales of $75$ m for the former and $1.5$ km for the latter, about $20$ long waves pass through a given radar cell, so that the mean surface level within that patch oscillates in phase with the long wave with a damping factor of about $1/20$. The amplitude of this damped oscillation remains large compared to the amplitude of the Bragg wavelength (because the ratio of the peak wavelength to the Bragg wavelength is $75/9.3\sim 8$ and the ratio of the corresponding amplitude is $\sim 8^{2}$ for a $k^{-4}$ wave spectrum). In addition, there are also wave groups of comparable or larger size to the radar cell. These groups are responsible for very low frequency modulation of the radar signal.
Since both the peak wave frequency $\omega_{p}$ and the current-induced frequency shift $\omega_{u}$ are much smaller than the Bragg frequency $\omega_{B}$, we can rewrite eq. (5) as the sum of an analytical signal $s_{a}$ (respectively, anti-analytical signal $s_{aa}$) that contains only the positive frequency side (respectively, negative frequency side) of the Bragg peaks:
$s(t)=s_{a}(t)+s_{aa}(t)=A^{+}(t)e^{i\omega_{B}t}+A^{-}(t)e^{-i\omega_{B}t},$ (7)
where
$A^{+}(t)=(R^{+}a^{+}e^{i\omega_{p}t}+R^{-}a^{+}e^{-i\omega_{p}t})e^{i\omega_{u}t}$ (8)
and
$A^{-}(t)=(R^{+}a^{-}e^{i\omega_{p}t}+R^{-}a^{-}e^{-i\omega_{p}t})e^{i\omega_{u}t}$ (9)
Assuming the complex amplitudes $R^{\pm}$ and $a^{\pm}$ to be uncorrelated due to random phase shifts between approaching and receding waves, we obtain that the variance of the analytic signal combined with the variance of the anti-analytic signal correlates with the variance of the ocean’s free surface associated with long waves we have:
$\mathrm{\langle\lvert A^{+}\rvert^{2}\rangle}+\mathrm{\langle\lvert A^{-}\rvert^{2}\rangle}\sim\mathrm{\langle\lvert R^{+}\rvert^{2}\rangle}+\mathrm{\langle\lvert R^{-}\rvert^{2}\rangle},$ (10)
Now from eqs (8) and (9) it appears that the absolute amplitudes $\lvert A^{\pm}\rvert$ can be extracted from the complex radar signal by taking the absolute value of its analytical and anti-analytical components:
$\lvert A^{+}(t)\rvert=\lvert s_{a}(t)\rvert,\ \lvert A^{-}(t)\rvert=\lvert s_{aa}(t)\rvert$ (11)
By construction the variables $\lvert s_{a}\rvert$, $\lvert s_{aa}\rvert$ and $\lvert R^{\pm}\rvert$ can be assumed to be Rayleigh-distributed random variables (as the modulus of complex Gaussian random variables). For this type of variable, the variance is a constant proportion of the second moment (i.e. $\mathrm{Var}{(X)}=\alpha\mathrm{\langle\lvert X\rvert^{2}\rangle}$ for some $\alpha>0$, here $\alpha=(1-\pi/4)$). It follows from (eq. 6) that the $H_{s}$ can be obtained as:
$H_{s}^{2}=C(\langle\mathrm{Var}{(\lvert s_{a}\rvert)}\rangle+\langle\mathrm{Var}{(\lvert s_{aa}\rvert)}\rangle)$ (12)
where $C$ is some calibration factor that must be set for each range and azimuth. This factor is calculated as the proportional coefficient between the mean value for the whole $2020$ year of the buoy $H_{s}$ and the mean value of the sum of $\lvert A^{+}\rvert$ and $\lvert A^{-}\rvert$ variances for the same period.
Figure 5: a) Real part of the received radar signal time series for the range $10$ and azimuth $75$ after beamforming (black) with the modulus of its analytical and anti-analytical signal (red) for September $27$, $2020$, at $23$h where $H_{s}$ and $T_{p}$ from CANDHIS are $2.93\ m$ m and $7.5\ s$, respectively. b) Modulus of the analytical signal for the same time. c) Power spectrum from the sum of the analytical and anti-analytical signals ($\sigma_{\lvert sa\rvert}(\omega)+\sigma_{\lvert saa\rvert}(\omega)$) for the same time. The common wave frequencies in the Mediterranean Sea are marked with a red rectangle where the maximum power spectrum is related to the peak frequency ($\omega_{p}$)
Since the amplitude modulation $\lvert A^{\pm}(t)\rvert$ follows the variation of long waves, the peak frequency ($\omega_{p}$) can be identified as the dominant frequency in the associated power spectrum:
$\omega_{p}=\mathrm{Argmax}(\sigma_{\lvert sa\rvert}(\omega)+\sigma_{\lvert saa\rvert}(\omega)),$ (13)
where $\sigma_{\lvert sa\rvert}$ and $\sigma_{\lvert saa\rvert}$ are the power spectra associated with the absolute value of the analytical and anti-analytical radar signals respectively. The $T_{p}$ can be derived from the relationship $T_{p}=1/\omega_{p}$. Figure 5 a) shows an example of the complex radar signal time series recorded on the antenna receiving array after beamforming processing. A slow amplitude modulation is visible, marked by the red lines. Part of this modulation is due to the beat frequency caused by the shift of the surface current from the Bragg frequency (see Guérin and Grilli, (2018)) and another part is due to the long wave groups (eqs. (8) and (9) ). This last term can be isolated by taking the modulus of the analytical/anti-analytical signal (see figure 5, b)), which has the effect of canceling out the Bragg component and the current-induced frequency shift. The peak wave frequency is sought at the maximum of the corresponding power spectrum (Figure 5, c)) in the range $0.1$ to $0.4$ Hz, which contains the dominant wave frequencies for the Mediterranean Sea. Only the positive part of the power spectrum is shown as it is symmetric.
4 Results and Discussion
This section is divided into the estimation of $H_{s}$ on the one hand and $T_{p}$ on the other since the precision of the estimation of these two variables is quite different. In this way, the problems and limitations encountered in the estimation of each variable can be discussed separately.
4.1 Significant wave height Estimation
The results for $H_{s}$ for the whole year $2020$, obtained using data from the Fort Peyras station only, are shown in Figure 6. Only minimal post-processing has been done by removing the obvious outliers $H_{s}>4$ m and $T_{p}>10$ s in the hourly estimates and performing a moving median average over 3 consecutive range bins (i.e. 4.5 km) and 5 consecutive azimuths (i.e. 5 degrees) to smooth the data and fill in some gaps. These results are shown along the azimuth $40$, as defined in Figure 1, at different distances from the radar station: in the short range (radar cell $8$ at $12$ km, hereafter referred to as $r8$); in the medium range (radar cell $20$ at $30$ km, $r20$); in the long-range (radar cell $30$ at $52.5$ km, $r30$). As can be seen, the results obtained using the SODS method are in good agreement with the data collected by the CANDHIS buoy. However, the accuracy of this method decreases with increasing distance from the radar. When we reach a distance of $20$ ($30$ km from the radar), it is no longer possible to estimate $H_{s}$ because the second-order level of the energy spectrum becomes indistinguishable from the noise. For this reason, Figure 6, c (in the range $35$ to $52.5$ km from the radar) does not show the results obtained with the SODS method. On the contrary, the estimate based on the FOBP is still relevant, if not accurate, at a greater distance.
Figure 6: Time series of $H_{s}$ (m) along the azimuth $40$ at a) range $8$ ($12$km), b) range $20$ ($30$km) and c) range $35$ ($52.5$km). The black line represents the data from the CANDHIS buoy located at Porquerolles, the grey line is the results of the WaveWatch III model run in the area, the red and blue lines correspond to the $H_{s}$ obtained using the SODS and the FOBP method from the radar data, respectively.
Table 1: Statistical performances of the SODS and FOBP methods measured over the year $2020$ along the central azimuth $a40$ at cell range $8$, $20$, and $35$ ($12$, $30$, and $52.5$ km, respectively) and compared with the ground truth of the CANDHIS buoy. The first index is the median-product correlation coefficient ($R^{\ast}$), the second index is the RMS difference (RMSE) and the last index is the difference of variances ($\Delta V$) in the time series.
$R^{\ast}$ RMSE ($m$ or $s$) $\Delta V$ ($m^{2}$ or $s^{2}$)
SODS FOBP SODS FOBP SODS FOBP
$H_{s}$ r8 $0.92$ $0.77$ $0.39$ $0.66$ $0.002$ $0.15$
r20 $0.88$ $0.72$ $0.66$ $0.61$ $0.33$ $0.06$
r35 $0.59$ $0.64$ $0.03$
$T_{p}$ r8 $0.56$ $0.69$ $1.60$ $1.67$ $1.71$ $0.11$
r20 $0.65$ $0.66$ $1.39$ $1.57$ $1.43$ $0.43$
r35 $0.46$ $1.73$ $0.92$
In order to compare both methods more extensively, the $H_{s}$ scatterplots for the $r8$ and $r20$ ranges (those for which SODS and FOBP results are available) along the $40$ azimuth are shown in Figure 7. It can be seen how for the SODS method the estimates are less scattered, although it overestimates the value of $H_{s}$ in all its value ranges, being this more evident farther away from the radar ($r20$). On the other hand, although the FOBP method shows a larger spreading, the observed deviation from the reference values is more symmetrical, neither overestimating nor underestimating the buoy $H_{s}$ value. It is also worth noting that the $H_{s}$ estimation with the FOBP method is more stable as we move away from the radar; we can see how for $r8$ and $r20$ similar shapes are obtained.
Figure 7: Scatter plot of radar $H_{s}$ from (a and c) SODS method and (b and d) FOBP method versus $H_{s}$ from CANDHIS buoy where each point is colored by the spatial density of nearby points. The spatial locations are along the $40$ azimuth and the ranges $8$ (a and b) first row, and $20$ (c and d) second row.
The respective performances of the $2$ methods have been quantified with a few statistical indices, using the CANDHIS buoy data as ground truth. The first index is the median product correlation coefficient $R^{\ast}$ (Shevlyakov, (1997)), which is more robust than Pearson’s correlation coefficient in series with noise and jumps. This index has already been used in the context of HF radars to assess surface currents (Dussol et al., (2022)) and we found it useful in the context of wave estimation. It is defined as
$R^{\ast}=\frac{\mathrm{Med}^{2}(\lvert a\rvert)-\mathrm{Med}^{2}(\lvert b\rvert)}{\mathrm{Med}^{2}(\lvert a\rvert)+\mathrm{Med}^{2}(\lvert b\rvert)}$ (14)
where:
$\begin{split}a&=H_{s\;radar}-\mathrm{Med}(H_{s\;radar})+\left(H_{s\;CANDHIS}-\mathrm{Med}(H_{s\;CANDHIS})\right);\\ b&=H_{s\;radar}-\mathrm{Med}(H_{s\;radar})-\left(H_{s\;CANDHIS}-\mathrm{Med}(H_{s\;CANDHIS})\right);\end{split}$ (15)
The second index is the Root Mean Square Error (RMSE) between the directional $H_{s}$ data from the radar and those from the buoy, which are taken as reference values.
Note that the RMSE is not a perfect indicator of accuracy as it can be artificially increased by the presence of outliers or, conversely, kept artificially low by taking a constant median of the time series. It is therefore useful to complement this quantity with an additional statistical index that compares the variability of the respective time series. This can be done simply by evaluating the difference $\Delta V$ of their respective variances. The results for the $2$ methods and the $3$ statistical indices are summarised in table 1 for the locations shown in figure 6 ($a40$ and $r8$, $r20$ and $r35$). In the short and medium range, the SODS estimate remains highly correlated with the buoy measurement ($R^{\ast}\simeq 0.9$), but its RMSE deteriorates from $0.39$ m to $0.66$ m, as does the difference in variances ($\Delta V$), which increases by two orders of magnitude; this difference is visually illustrated in Figure 6, b). On the contrary, the performance of the FOBP, although not optimal, remains remarkably stable with increasing range ($R^{\ast}\simeq 0.6-0.8$ and RMSE0.65similar-to-or-equalsabsent0.65$\simeq$ 0.65≃ 0.65 m). In this case, the $\Delta V$ remains more constant and even decreases as we move away from the radar. Note that the obtained RMSE values should be compared with the mean value of $H_{s}$ observed in the Toulon region (Figure 1a), which ranges from about 0.9 m to 1.2 m with increasing distance, and with the intrinsic RMS difference between the $H_{s}$ at the radar and buoy location, which ranges from $0.1$ m to $0.4$ m (Figure 1b).
A spatial distribution of these statistics in the Toulon radar area is shown in figure 8. It can be seen that the above statements are true not only along the azimuth $40$ but over the whole radar coverage. It should be noted that the SODS method is very sensitive to any electromagnetic interference that may be present in the area, and for this reason, the results obtained in the shadow of the Giens tombolo (east of the radar coverage) are of poorer quality. On the other hand, the FOBP method, based on the first-order Bragg signal, is less affected by such interference and remains satisfactory even in the vicinity of the Giens peninsula. However, the latter method suffers from another type of interference, namely the power grid artifact discussed in section 3.1, which distorts the received signal at distances of about $37.5$ and $45$ km from the radar site. These data have been excluded from the results.
Figure 8: Spatial maps of the $3$ statistical performance indicators for the radar measurement of $H_{s}$ obtained for the year $2020$ with the SODS (left column, a, c, e) and FOBP (second column, b, d, f) methods: 1) $R^{\ast}$ (a, b); 2) RMSE (c, d); 3) $\Delta V$ (e, f). The black arcs mark the ranges r8, r20, and r35 ($12$, $30$, and $52.5$ km) respectively, and the black radial indicates the azimuth used for the calculations in table 1. The SODS is limited to r35, while the FOBP extends well beyond. The white circular bands in the maps in the right column correspond to the few areas where the low-frequency component of the radar signal is distorted by a power grid artifact.
4.2 Peak wave period Estimation
Regarding the results obtained for $T_{p}$ with the classical SODS method, we can observe in Figure 9 and Table 1 worse performances, with a reduced correlation coefficient ($R^{\ast}\simeq 0.6$) and a relatively large RMS error (RMSE1.5similar-to-or-equalsabsent1.5$\simeq$ 1.5≃ 1.5 seconds, that is about $24\%$ of the mean peak period $6.14$ seconds). From Figure 9 it can be seen that the SODS radar estimate is not able to follow the full variability of $T_{p}$, especially towards the lowest values. This could be due to the reduced accuracy of the asymptotic formula (eq. 4) at low wind speeds, as suggested by recent numerical simulations (Guérin,, 2024). The $T_{p}$ estimated with the FOBP shows similar performances to SODS in terms of correlation coefficient and RMSE, even in the near range (Table 1). However, as can be seen from the plots in Figure 9, this new estimate has a much better ability to follow the fluctuations of $T_{p}$, even at its lowest values, and it is robust with increasing range. This is reflected in the difference in the variances of the time series, which is much smaller than with the SODS method (Table 1).
Figure 9: Time series of $T_{p}$ (s) along the azimuth $40$ at a) range $8$ ($12$km), b) range $20$ ($30$km) and c) range $35$ ($52.5$km). The black line represents the data from the CANDHIS buoy located at Porquerolles, the grey line is the results of the WaveWatch III model run in the area, the red and blue lines correspond to the $T_{p}$ obtained using the SODS and the FOBP method from the radar data.
The scatter plot analysis confirms these results (Figure 10). It is shown how the SODS method cannot capture $T_{p}$ less than $4$ s due to the limitation of distinguishing the second order of the spectrum from the noise, while the FOBP method, which does not have this limitation, can capture periods up to $2.5$ s. For $T_{p}$ the same happens as for the $H_{s}$ estimation, the SODS method, despite having a narrower shape, overestimates the buoy values, although in this case no deterioration is observed as we move away from the radar. As for the FOBP method, the shape obtained is symmetrical for high periods, while in the lower period range, this method underestimates the value of the buoy period. This may be due to the large number of wave periods of similar energy that can be found in a calm sea. In this situation, the FOBP method may give erroneous values, but close to the ground truth. This fact can be better explained in figure 5, c), where we can see that there are at least $3$ peaks at frequencies lower than $\omega_{p}$ with energy very close to that selected by the FOBP method. On the other hand, when high periods are found, we are usually in storm conditions where the waves are more extreme and more ordered in wavefronts, identified by a strong peak energy around $\omega_{p}$, making it easier to identify this value. For this reason, and although FOBP can solve for periods smaller than $4$s, this method is still at an early stage of development and may improve the estimation of $T_{p}$ at low period values in future work.
Figure 10: Same as Figure 7 for $T_{p}$.
The results highlighted with some radar cells ($r8$, $r20$, and $r35$) along the central azimuth ($a40$) have been confirmed with other radar cells (see figure 11). As for the $H_{s}$ estimation, the ranges are truncated at $30$ km for the SODS method, while for the FOBP method, the valid ranges extend up to $52.5$ km. The missing values around $19.5-24$km and $37.5-45$km (indicated by the white circle in figure 8.b, 8.d and 8.f) are those where the electrical artifact visible in figure 4 compromises the low-frequency content of the radar signal. In these maps we can see that for locations further away from $r8$ the correlation with the buoy is much better for the FOBP method than for the SODS method, being able to reach $r25$ with acceptable correlation and RMSE values (see figure 11.b and d). Furthermore, as mentioned above, the interpretation of the RMSE value for the SODS method requires some caution, since it is true that in some locations this statistic is better than for the FOBP method, but this is simply because the SODS method does not capture the variability of the $T_{p}$ signal (Figure 9), being unable to estimate low period values. For this reason, this analysis must be accompanied by the variance difference analysis (Figure 11, e) to understand the RMSE results as reliably as possible.
Figure 11: Same as Figure 8 for $T_{p}$.
4.3 Omnidirectional estimation
A challenging issue is the ability to estimate sea state parameters when the radar time series or Doppler spectra cannot be resolved in azimuth, as is the case for example with compact HF radars. To test this possibility, we applied the SODS and FOBP methods to the different antenna signals without any azimuthal processing. Omnidirectional Doppler spectra and omnidirectional power spectra of amplitude modulation (as defined in section 3.2) were obtained by processing each antenna channel individually and averaging over the $12$ antennas. Due to the variation of the radial flow over the radar coverage, this has the effect of broadening the Doppler spectrum compared to azimuthally resolved Doppler spectra. However, for moderate values of the radial current (which is the case in the Toulon area), this does not cause a critical overlap of the first and second-order regions in the omnidirectional Doppler spectra and the SODS can therefore be applied in the same way. The estimation of $H_{s}$ and $T_{p}$ over the year $2020$ was done in a similar way and the updated statistical power indices are shown in table 2. The evolution of the RMSE of these estimates as a function of the sole range is shown in Figure 12 and the directional estimate at azimuth $40$ and the RMSE from the CANDHIS buoy are given for comparison. It can be seen that the performance of the SODS method is very similar in the directional and omnidirectional cases, with only a slight deterioration in the $H_{s}$ estimate that becomes more pronounced as we move away from the radar location. On the contrary, there is a slight improvement for the $T_{p}$ estimate. This is related to the weak variation of these parameters over the radar coverage, as seen for the WW3 model (Figure 2). The same conclusion applies to the FOBP method for the estimation of $H_{s}$, with a very stable RMSE of about $0.6-0.7$ metres over the range from $8$ to $35$ (i.e. from $12$ to $52.5$ km). Although this RMSE is larger than that obtained with SODS in the short range (about $0.4$ m), it is already better than the latter method in the medium range from $13$ ($19$ km). This is not the case for $T_{p}$, which is very poorly estimated. This could be due to the smearing of the peak frequency of the amplitude modulation $A^{\pm}(t)$ when the radar cell is no longer resolved in azimuth.
Table 2: Same as Table 1 for the omnidirectional SODS and FOBP methods.
$R^{\ast}$ RMSE ($m$ or $s$) $\Delta V$ ($m^{2}$ or $s^{2}$)
SODS FOBP SODS FOBP SODS FOBP
$H_{s}$ r8 $0.90$ $0.73$ $0.44$ $0.60$ $0.058$ $0.085$
r20 $0.82$ $0.64$ $0.83$ $0.63$ $0.408$ $0.117$
r35 $0.42$ $0.71$ $0.010$
$T_{p}$ r8 $0.58$ $0.35$ $1.30$ $1.95$ $1.667$ $1.548$
r20 $0.52$ $0.09$ $1.40$ $2.51$ $1.572$ $2.004$
r35 $0.04$ $2.86$ $2.435$
Figure 12: RMSE of directional (solid lines) and omnidirectional (dashed lines) a) $H_{s}$ in $m$ and b) $T_{p}$ in $s$ from the SODS method (red line) and FOBP method (blue line). The dashed black line corresponds to the mean annual value of $H_{s}$ and $T_{p}$ in the radar coverage. The directional estimation is taken along the azimuth a40.
The scatter-plot of $H_{s}$ estimates from omnidirectional radar measurement is shown in Figure 13 for the same two ranges as considered previously ($r8$ and $r20$) and confirms the robustness of this estimation when no prior selection of bearing can be accomplished, finding the same characteristics as in the directional estimation of each of the methods (see discussion in Section 4.1). It becomes even more evident here that the SODS method overestimates $H_{s}$ as we move further away from the radar.
Figure 13: Scatter plot of omnidirectional radar $H_{s}$ from (a and c) SODS method and (b and d) FOBP method versus $H_{s}$ from CANDHIS buoy where each point is colored by the spatial density of nearby points. The spatial locations are along the ranges $8$ (a and b) first row, and $20$ (c and d) second row.
4.4 Comparison with the literature
A number of studies have been published in the literature on the estimation of wave heights and periods from a single HFR station and a single radar frequency. Some of them reported systematic comparisons with in situ instruments over a significantly long period (at least several months) using either phased array (Wyatt et al.,, 2003; Ramos et al.,, 2009; Gomez et al.,, 2015; Lopez et al.,, 2016; Lopez and Conley,, 2019; Alattabi et al.,, 2019; Al-Attabi et al.,, 2021) or compact radar systems (Roarty et al.,, 2015; Zhou and Wen,, 2015; Long et al.,, 2011; Atan et al.,, 2016; Saviano et al.,, 2020; Basañez et al.,, 2020). All of these methods are based on the exploitation of the second-order Doppler spectrum using either Barrick’s original methods, the full wave spectrum inversion method, or empirical methods developed by the radar manufacturers. The performance obtained varies considerably depending on the geographical location, the radar frequency, the measurement range, the antenna array, the inversion method chosen, the data post-processing choices, etc. However, two general trends can be identified within this diversity of results. First, the estimation of $H_{s}$ with phased array systems is generally slightly better than that with compact systems, with high Pierson correlation coefficients ($>0.9$) and RMSE often in the order of 30-40 cm and less than about 60 cm in all studies. The observed RMSE with compact systems are generally larger, in the order of 40-80 cm, with a large scatter of values and in some cases very large RMSE ($>1$ m); the associated correlation coefficients are generally smaller, mostly in the range of 0.5-0.9, except for the study Atan et al., (2016), which reports exceptionally good performances of 18-29 cm RMSE and 0.78-0.86 Pierson correlation coefficient. However, the decision to eliminate the $>25$% outliers (when compared with buoy measurement) before calculating the statistics is questionable. The second point is that the estimation of $T_{p}$ is less systematic and generally of lower quality than that of $H_{s}$, with RMSE of the order of $1-3$ seconds (to be compared with the usual range of periods, say $6-10$ seconds) and Pierson correlation coefficients of the order of 0.5-0.7.
Our results are in agreement with these trends. The RMSE obtained for the estimation of $H_{s}$ with SODS ranges from $0.39$ to about $0.5$ m for a wide angular sector in the near and medium range (up to about 20 km, see Figure 8c) and is slightly larger in the omnidirectional case, which is more representative of a compact antenna array. The estimation of $T_{p}$ is much coarser, with a lower correlation coefficient and an RMSE of at best $1.3$ seconds (that is a relative error of the order of $25$%). Note that we could have improved the statistical comparisons by additional post-processing of the data (elimination of outliers, jumps, etc.), but we chose to limit the impact of post-processing, which can be misleading in the evaluation of the method.
5 Conclusion
In this work, we have tested the accuracy of the classical Barrick’s method for estimating the main sea state parameters based on weighted integrals of the second-order backscattered Doppler spectrum. For this purpose, we analyzed a one-year HFR data set from a monostatic station in the Toulon region. In this region, the wave characteristics are homogeneous, allowing us to use both the directional and omnidirectional signals to estimate the significant wave height and peak period within the radar coverage area. The results presented in this study have been compared with the data collected by the CANDHIS buoy located near the island of Porquerolles. These data from the buoy have been considered as the ground truth for the analyses presented in this study. For the classical Barrick’s method with directional radar signal, the results obtained for significant wave height have a good accuracy compared to the time series of the CANDHIS buoy, with a very high correlation and a low RMSE value between both of them (close to $90\%$ and around $0.4$m, respectively, up to $30$km from the radar). Regarding the estimation of the wave peak period, this method shows some deficiencies in capturing low period values. This may be due to the limitation of differentiating the second-order spectrum from the signal noise or to a lower precision in the asymptotic formula of the weighting function at low wind speeds. The main limitations of this method are its sensitivity to electromagnetic interference and its degradation as we move further from the radar, reaching a point, in this study about $30$ km from the radar, where the results are unreliable.
In addition, a new method based on the modulation of the first order of the backscattered Doppler spectrum was presented in this study. This method allows wave characteristics to be estimated at greater distances from the radar since the first order of the spectrum is generally much more energetic than the second order and is easier to distinguish at great distances. The results obtained for the same year selected for Barrick’s method using the directional signal of the radar show that this new method gives a good quality in the estimation of the significant wave height, obtaining a better coverage in the estimation (up to 52.5 km compared to 30 km for the Barrick’s method). As for the estimation of the peak period, this new method gives better results than Barrick’s method even at distances close to the radar, with correlations close to $80\%$ and an RMSE of $1.5$s. While this method gives good results in estimating the peak period of the wave, it underestimates the buoy values for small periods, which requires some improvement in future work. The main limitation of this new method so far is that, at least for the $H_{s}$, it requires a preliminary calibration procedure with some ground truth measurement. This is not the case for the SODS method, which is based on the ratio of the first and second-order Doppler spectra and is therefore insensitive to absolute calibration.
In this study, in addition to treating the directional radar signal to estimate the wave characteristics, omnidirectional signals have also been evaluated to test the quality of both methods when it is impossible to perform the beamforming technique and to distinguish from which azimuth the signal arrives. It was concluded that both methods can still give good results for estimating the significant wave height but not only Barrick’s method remains relevant for the estimation of the peak wave period in the omnidirectional case. The first trials of the FOBP method were presented in this paper with promising results. The technique will be pursued and hopefully improved in future works.
Acknowledgment
Part of this research and the first author (VMM) have been funded by the Agence Nationale de la Recherche (ANR) under grant ANR-22-ASTR-0006-01 (ROSMED project: “Radar à Ondes de Surface en MEDiterranée”). We acknowledge the MOOSE program (Mediterranean Ocean Observing System for the Environment) coordinated by CNRS-INSU and the Research Infrastructure ILICO (CNRS-IFREMER). We thank the Parc National de Port-Cros (PNPC), the “Association Syndicale des Propriétaires du Cap Bénat” (ASPCB) as well as the Group Military Conservation (GMC) and the Marine Nationale for hosting our radar installation in Porquerolles, Cap Bénat and Fort Peyras, respectively. In addition, we would like to thank Alejandro Cáceres-Euse for providing the WWIII simulation results in the study region, which were used as a complementary dataset to validate the methodology here presented.